For each of the following independent cases (A-E), compute the missing values in the table: Total Manufacturing Cost CaseCost Prime Conversion Direct Direct Manufacturing Cost Materials Labor Overhead A $ 9,440 $ 16,450 3,990 $11,000 10,1507,480 42.400 20,500 11,450 26,000 43,630 56,200 110,700 49,250 69,750 13,700 24,700 57,300
Answer #1

Answers

  • Missing amounts are calculated by using following equations:
    --Direct Material + Direct Labor = Prime Cost
    --Direct Labor + Manufacturing Overhead= Conversion Cost
    --Direct material + Direct labor + manufacturing overhead = Total manufacturing cost.
  • Requirement

CONVERSION COST

PRIME COST

Case

Prime Cost

Conversion Cost

Direct Materials

Direct Labor

Manufacturing Overhead

Total Manufacturing Cost

A

$                                        9,440

$                16,450

$                 3,990

$                 5,450

$               11,000

$              20,440

B

$                                      17,630

$                33,480

$               10,150

$                 7,480

$               26,000

$              43,630

C

$                                      56,200

$              110,700

$               42,400

$               13,800

$               96,900

$            153,100

D

$                                      31,950

$                49,250

$               20,500

$               11,450

$               37,800

$              69,750

E

$                                      32,600

$                38,400

$               18,900

$               13,700

$               24,700

$              57,300

--Working

CONVERSION COST

PRIME COST

Case

Prime Cost

Conversion Cost

Direct Materials

Direct Labor

Manufacturing Overhead

Total Manufacturing Cost

A

9440

16450

3990

=16450-11000

11000

=3990+5450+11000

B

=10150+7480

=7480+26000

10150

7480

26000

43630

C

56200

110700

42400

=56200-42400

=110700-13800

=42400+13800+96900

D

=20500+11450

49250

20500

11450

=49250-11450

69750

E

=18900+13700

=13700+24700

=57300-13700-24700

13700

24700

57300